OpenStudy (anonymous):

I NEED HELP PLEASE I WILL GIVE MEDALS AND WHATEVER YOU NEED!!!! Part A: Divide (8x4y3 + 4x3y2 - 2x2y - 12x2y4) by -2x2y. Show your work, and justify each step. Part B: How would your answer in Part A be affected if the x2 variable in the denominator was just an x? Part C: What is the degree and classification of the polynomial you got in Part A?

OpenStudy (anonymous):

You can split this up and reduce separately if that makes it easier. \[\frac{ 8x^4y^3 }{ -2x^2y }+\frac{ 4x^3y^2 }{ -2x^2y }-\frac{ 2x^2y }{ -2x^2y }-\frac{12x^2y^4 }{ -2x^2y }\]

OpenStudy (anonymous):

i just have to complete that to get A?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

basically what @slobo_da_hobo is telling you is to cancel the number of x's and y's for example X^3/x2 = x * x * x/x* x = x (3x's divided by 2 x's gives one x)

OpenStudy (anonymous):

i'll do the first one: \[8x^4y^3/-2x^2y = -4x^2y^2\]